OpenStudy (anonymous):

Sorry accidently closed the last one. Simplify each expression. Use positive exponents. (x^–2 y^–4 x^3)^–2 (x^–2 x^3)^–2 (y^–4)^–2 (x^1)^–2 (y^8) Is that close to it? I am completely lost at math but not looking for a get an answer quick just can't figure this out and learn it. Any and all help would be wonderful. Thanks.

OpenStudy (anonymous):

ok so first one multiply the x's in the parenthesis together and you get x^3/x^2 which is x, do you see that?

OpenStudy (anonymous):

ok so you do that and you have (y^(-4) * x)^(-2) then you just raise each term in the parenthesis to the power of -2, since all the terms are multiplied, so you have x^(-2) * y^(-4)^(-2) which is x^(-2) * y^(8), just multiply the exponents by the power you are raising each term to which is -2. by the exponent rule (a^b)^c = a^(b*c) , then from x^(-2) * y^(8) you get rid of negative exponents by moving them to denominator and you have (y^8) / (x^2)

OpenStudy (anonymous):

see if you can figure out the rest from that example

OpenStudy (anonymous):

Sorry about late reply. I think I got it. \[\frac{ x^{4} y^{8} }{ x^{6} } = \frac{ y^{8} }{ x^{2} }\]